OpenStudy (anonymous):

Tom has 7 more silver dollars than jimmy. Together they have a total of 71 silver dollars. How many silver dollars does Tom have?

OpenStudy (pfenn1):

X=# of silver dollars Jimmy has Y=# of silver dollars Tom has X+Y=71 Y=X+7 X+Y=X+(X+7)=71 2X=64 X=32 Jimmy has 32 silver dollars and Tom has 39.

OpenStudy (anonymous):

Thanks, is that the final answer 39?

OpenStudy (pfenn1):

yes, we came up with X so plug that back into the equation Y=X+7=32+7=39